Much concern has been expressed regarding the practice of using nitrates as meat preservatives. In one study involving possible effects of these chemicals, bacteria cultures were grown in a medium containing nitrates. The rate of uptake of radio-labeled amino acid (in dpm, disintegrations per minute) was then determined for each culture, yielding the following observations.

7251 6871 9639 6866 9098 5844 8955 7972
7068 7498 7887 8172 7528 8728 7462

Suppose that it is known that the mean rate of uptake for cultures without nitrates is 8000.

Do the data suggest that the addition of nitrates results in a decrease in the mean rate of uptake? Test the appropriate hypotheses using a significance level of 0.10. (Use a statistical computer package to calculate the P-value. Round your test statistic to two decimal places and your P-value to three decimal places.)

t =

p-value =
Answer #1

Ans:

sample mean=7789.267

sample standard deviation,s=1004.054

n=15

Test statistic:

t=(7789.267-8000)/(1004.054/sqrt(15))

t=-0.81

df=15-1=14

p-value=tdist(0.813,14,1)=0.215